OpenStudy (anonymous):

Help please!!!! Find the exact zeroes of the function in the interval [0,2pi): f(x) = sin x/2 - cos x

OpenStudy (anonymous):

use the half angle formula...

OpenStudy (anonymous):

\[\sin \frac{x}{2} - cosx = 0\]

OpenStudy (anonymous):

\[\sin \frac{x}{2} = \pm \sqrt{\frac{1-cosx}{2}} \]

OpenStudy (anonymous):

now your equation can be written as:\[\pm \sqrt{\frac{1-cosx}{2}} = cosx\]

OpenStudy (anonymous):

squaring both sides gives: (1-cosx)/2 = cos^2(x) 2cos^2(x) + cosx -1 = 0 notice this is a trig equation in quadratic form so see if you can factor: (2cosx - 1)(cosx + 1) = 0 equate each factor to 0 and solve. ...should be easy from here.